# Social Norms in Cinema: A Cross-Cultural Analysis of Shame, Pride and Prejudice

### v4

@@ 4.2.1 Manual Evaluation
4.2.2 Thematic Analysis
Twenty-four clusters for shame-related reasons and fifteen for pride-related reaosns were formed using agglomerative clustering ( See Tables A13 and A14 for clustering parameters). The clusters were manually assigned a theme (as depicted on the Y-axis of Fig. 2 and 3) after analyzing the ten closest samples based on cosine distance within the cluster (See Table A11 and A12 for examples in each cluster). We computed the relative association for each cluster with Bollywood and Hollywood using eq. 1 and performed the Barnard-Exact Test Barnard (1947) with the Yates Correction Yates (1934) to test if the possibility of observing norms related to pre-assigned themes is statistically different across movie industries.
@@ 4.2.2 Thematic Analysis
Here, $\mathcal{D}_{industry_{t_{i}}}$ represents dialogues to $industry\in\{holly,bolly\}$ and ${t_{i}}\in themes$ (as depicted on Y-axis in Fig 2 and 3).
Figure 2: Relative association ($\Delta$) of Bollywood and Hollywood to themes obtained from agglomerative clustering performed on shame-related norms. * indicates significant difference i.e., $p>0.05$. See Table A15 for statistics and p-value.
@@ 4.2.3 How is shame/pride attributed across genders?
Figure 3: Relative association ($\Delta$) of Bollywood and Hollywood to themes obtained from agglomerative clustering performed on pride-related norms. * indicates significant difference i.e., $p>0.05$. See Table A16 for statistics and p-value.
Figure 4: Relative association ($\Delta$) of emotions pride and shame gender-wise in Hollywood and Bollywood dialogues. A higher positive score indicates a stronger association of gender with pride.
@@
Here, $\mathcal{D}_{pride_{g}}$ represents movie dialogues having pride-related words where the target of pride is $g\in\{male,female\}$, $\mathcal{D}_{shame_{g}}$ represents movie dialogues having shame-related word where the target of shame is $g\in\{male,female\}$ and $\mathcal{D}_{g}$ represents total movie dialogues grouped based on gender. As depicted in Fig. 4, Hollywood movies are pride-oriented, whereas Bollywood movies are shame-oriented. Females are attributed more shame, and the difference (male-female) in the expression of pride and shame is $0.16$ for Hollywood and $0.21$ for Bollywood.
Social expectations evoking shame are similar across movie industries (See Fig. 5 and 6). Sexuality (e.g., promiscuity, immodesty) is the dominant theme for women in both movie industries whereas males are shamed for incompetency in both industries. Women across movie industries express more pride in family roles. Men-pride in Bollywood is centered on justice, winning and bravery whereas it is duty, self-identity and winning in Hollywood (See Fig. A1 for gender differences in pride).
Figure 5: Gender differences in shame-related themes in Bollywood movies. * indicates significant difference i.e., $p>0.05$.
Figure 6: Gender differences in shame-related themes in Hollywood movies. * indicates significant difference i.e., $p>0.05$.
@@ 5 Discussion
We release a cross-cultural dataset of shame (6477 samples) and pride (3968 samples) related expressions with implicit reasons behind them. To the best of our knowledge, the only related corpus is the GoEmotions dataset Demszky et al. (2020) comprising 817 samples for embarrassment and 452 samples for pride. This corpus can serve the task of developing computational models for detecting shame/pride expressions and recognizing how appraisals of such emotions vary -- enabling culturally cognisant human-computer conversations for diverse users Kim et al. (2011) and cross-cultural intent translation [^5] DARPA Computational Cultural Understanding: https://www.darpa.mil/program/computational-cultural-understanding.
@@
Shame is a highly undesirable self-focused emotion emphasizing incompetency and failures and is rarely used in the U.S. Cohen (2003); Boiger et al. (2013) whereas Eastern cultures use shame more frequently and have a desirable affect Lim (2016). Notably, we also observe infrequent use of shame in Hollywood movies and self-focus (reflecting internal shame) (See Table 2 and 3 for frequency distribution and psychosocial constructs). Unlike in Western nations, shame is interdependent in collectivist communities Wong and Tsai (2007). The empirical analyses reveal other-focused (also, the public nature of shame) in Bollywood movies (See Table A7). Moreover, the contrasting tenses coupled with varying emotions, i.e., past + sadness in Hollywood vs. present + anger in Bollywood, reflect their varying goals, i.e., remorse for past failures/losses vs enforcing conformity Wong and Tsai (2007). Pride-related discourse in Hollywood is duty and achievement-focused, in line with prior findings underlining the increasing significance of "success" with growing individualism Cohen (2003), whereas in Bollywood, pride is centered around collective achievement (we, achievement, prosocial). It is also interesting to note linguistic markers such as determiner (used with objects/nouns) in Bollywood indicating more materialistic pride, compared to personal pronouns and social references in Hollywood reflecting more personal pride (See Tables A8, A10 for all significant categories correlated with pride).
@@
Eliciting reasons leading to shame and pride expressions revealed social expectations across cultures. We observed high specificity in extracted reasons revealing cultural subtleties (e.g., desire for a son in Bollywood vs. returning calls after date night in Hollywood). To tease apart ways in which cultures differ, we first considered mapping social expectations/generated reasons to Schwartz’s Theory of Values Schwartz (2012). However, social expectations were found to have contrasting values depending on the culture. Consider "refusing to marry", an instance of non-conformity in Indian society, whereas it is an instance of self-direction in a Western context. To effectively understand the cross-cultural variations in social expectations underlying shame/pride, we performed hierarchical clustering and empirically picked the distance after manually analyzing the quality and granularity of clusters. The situations connected to shame and pride in the U.S. society (i.e., lacking accountability, poverty, harm) vs Indian society (e.g., inappropriate sexual behavior, gender roles, betrayal - see Fig 2 and 3) align with known cultural tendencies (i.e., individualist vs collectivist) of both nations Triandis (1988).
@@
The vision of safe and accountable AI is centered on LLMs’ moral and value alignment. The dominant approach for norm discovery involves prompting LLMs, sometimes coupled with a verification step such as an entailment test or underlying emotion (negative emotion $\rightarrow$ norm violation) Jiang et al. (2021); Fung et al. (2022); CH-Wang et al. (2023). However, most of the social situations and the human annotators employed to label those situations reflect English beliefs and ethics. Our approach to extracting culture-specific social expectations and attitudes using social emotions overcomes this limitation. While the clustering was performed on the reasons extracted from movie conversations, the cluster themes are similar to patterns seen during LIWC analysis, validating our approach for overcoming Anglocentric bias in LLMs for norm discovery Ziems et al. (2022).
@@
Lastly, we examined how social expectations associated with shame/pride are attributed to men and women across cultures influencing their social behaviors; women are less assertive across cultures to avoid negative attribution Amanatullah and Morris (2010); Ferguson and Eyre (2000) and ask for less during negotiation than men Arnold and McAuliffe (2021). We demonstrated that shame, a negative self-conscious emotion expressing devaluation, is targeted towards women more than men, whereas pride, endorsing increased social value, is used for more men than women in movies. The similarity in social expectations (e.g. regulating sexual behaviors in women vs shaming incompetency in men) across cultures is surprising (See Fig 5 and 6). It is important to characterize the source and acquired social expectations before their use for achieving alignment in LLMs otherwise we risk social bias.
@@ 6 Conclusion
We introduce a multi-cultural dataset of shame/pride-related expressions and the underlying social expectations. This study demonstrates (a) cross-cultural linguistic differences in shame and pride-related expression, offering insights into their functions across cultures, (b) cultural dichotomy in social expectations, and (c) more "social sanctions" and fewer "endorsements" to the female gender in movies. Future work can utilize our dataset for culturally aligned LLMs and build social emotion perception and appraisal in human-computer interactions. Additionally, our analysis of shame/pride is the first of its kind; we hope future NLP researchers will build upon this work to investigate social expectations in LLMs from a multicultural, social emotion-based perspective.
@@
Social norms discovery is a crucial component in programs [^6] ALIGN-https://www.alignplatform.org/learning-collaborative designed for social and behavioral change to promote equity, social justice, and well-being Mauduy et al. (2022); Bonan et al. (2020). Further work explores style as a product of norm differences Havaldar et al. (2023a). Social psychology investigates social norms (descriptive vs injunctive) to design experiments for understanding behaviors such as self-regulation, persuasion Cialdini et al. (1990) and decision-making Gavrilets (2020); Bhanot (2021) to promote collective-level change in societies.
@@
Social norms mutate as society evolves. We acknowledge that our dataset of movies (released post-1990) may reflect social norms that are less characteristic of contemporary society. Moreover, countries like India and America contain a mix of cultures. The captured norms may not reflect the cultural variations, for example, between regions (e.g., East Coast vs West Coast in the U.S.A or North India vs South India). Movies also exaggeratedly depict the world around us (e.g., weddings, criminal activities, sexual abuse, etc.), and we caution against stereotyping cultures based on movie-based norms.
The dominant language in Bollywood movies is Hindi and our analysis is based on their English translations which may not always be accurate, especially when the discourse is about concepts native to a culture. Relatedly, the LIWC may not have high coverage for such concepts. We did not compare the movie genre and acknowledge that situational/unrealistic norms (e.g., science fiction, comedy, etc.) could exist. Regardless, social norms associated with shame and pride are still relatable and reflect the target audience’s beliefs. It is also worth noting that gender norms in movies may not fully reflect the norms of today’s society. This study is conducted over a thirty-year long period and thus, the aggregated norms may not reflect the current trends.
@@ Acknowledgements
We acknowledge the use of ChatGPT, an AI language model developed by OpenAI, for its assistance in refining this manuscript.
@@ A.2 Examining Implicit shame
We prompted GPT-4 chat to identify (a) if a given discourse expresses shame, and (b) identify the behavior or action behind shame, for a randomly sampled set of 100 samples without explicit mentions of shame/pride each from Hollywood and Bollywood. We asked two human annotators to evaluate the correctness of GPT-4 chat-generated outputs.
@@ A.3 Prompts
The prompts designed to extract the subject experiencing self-conscious emotions, their gender, and the reason behind the emotions are provided in Tables A2 and A3.
@@
"You are an experienced social scientist and you study Indian society. Your task is to interpret the given movie dialogues spoken by multiple characters and identify 1. Who is being shamed? Predict the gender. If gender is not explicit, use the pronouns, name, and salutations to guess gender. 2. Identify the primary reason for shaming The answer should be short and in CSV format. $<$ shamed, gender, reason$>$ Input: And should we bow before others begging....them to marry our daughters? This shall not happen. Neither will the girls be alive here nor shall....we be ashamed of ourselves. You cannot kill the life which God has given. I won’t let you commit the sin. Output: girl’s parent, NA, not able to marry off their daughters Input: Black marketers are now in the open. And the thieves too Politics is in a great mess Shame on this system. There’s no democracy Get rid of these politicians The gong has struck..’’Our hearts are swaying to it’s beats’’ Output: System, NA, poor law and regulations"
@@
"You are an experienced social scientist and you study American societies. Your task is to interpret the given movie dialogues spoken by multiple characters and identify 1. Who is being shamed? Predict the gender. If gender is not explicit, use the pronouns, name, and salutations to guess gender. 2. Identify the primary reason for shaming. The answer should be short and in CSV format. $<$ shamed, gender, reason, prevalence $>$ Input: You still owe me 100. Remember? You stiffed Donny for 100 bucks? Cheapskate. Shame on you. Pay this man his C-note. Now I know why they call you the Snake. Output: NA, male, not returning borrowed money Input: You prey on your own people. You steal from your own people. Have you no shame!? - Huh? - Well, we’re still here. Man: Mr. Markopolos, it’s all yours. Output: Snake, male, stealing and preying on people"
@@ A.4 LIWC Correlation Results
Tables A7, A8, A9 and A10 contain the significantly correlated ($p<0.05$) LIWC categories, the most frequent five words for each category, pearson r and 95% confidence interval.
@@
YOU (you, your, you’re, yourself, you’ve) 0.161 0.000 [0.145, 0.178]
@@
AUXVERB (is, are, have, be, don’t) 0.029 0.001 [0.013, 0.046]
@@
Table A7: Psychosocial categories significantly correlated (p<0.05) with shame in Bollywood dialogues. p-values were corrected using Benjamini-Hochberg correction. The categories are arranged in decreasing order of correlation.
@@
Table A8: Psychosocial categories significantly correlated (p<0.05) with pride in Bollywood dialogues. p-values were corrected using Benjamini-Hochberg correction. The categories are arranged in decreasing order of correlation.
@@
Table A9: Psychosocial categories significantly correlated (p<0.05) with shame in Hollywood dialogues. p-values were corrected using Benjamini-Hochberg correction. The categories are arranged in decreasing order of correlation.
@@
Table A10: Psychosocial categories significantly correlated (p<0.05) with pride in Hollywood dialogues with 95% confidence intervals. p-values were corrected using Benjamini-Hochberg correction. The categories are arranged in decreasing order of correlation.

### v5

@@ 4.2.2 Thematic Analysis
Here, $\mathcal{D}_{industry_{t_{i}}}$ represents dialogues to $industry\in\{holly,bolly\}$ and ${t_{i}}\in themes$ (as depicted on Y-axis in Fig 4 and 5).
Figure 4: Relative association ($\Delta$) of Bollywood and Hollywood to themes obtained from agglomerative clustering performed on shame-related norms. * indicates significant difference i.e., $p<0.05$. See Table A15 for statistics and p-value.
@@ 4.2.3 How is shame/pride attributed across genders?
Figure 5: Relative association ($\Delta$) of Bollywood and Hollywood to themes obtained from agglomerative clustering performed on pride-related norms. * indicates significant difference i.e., $p<0.05$. See Table A16 for statistics and p-value.
Figure 6: Relative association ($\Delta$) of social emotions attributed to male and female. A higher positive score indicates a stronger association of gender with pride.
@@
Here, $\mathcal{D}_{pride_{g}}$ represents movie dialogues having pride-related words where the target of pride is $g\in\{male,female\}$, $\mathcal{D}_{shame_{g}}$ represents movie dialogues having shame-related word where the target of shame is $g\in\{male,female\}$ and $\mathcal{D}_{g}$ represents the number of movie dialogues spoken by $g$. As depicted in Fig. 6, Hollywood movies are pride-oriented, whereas Bollywood movies are shame-oriented. Females are attributed more shame, and the difference (male-female) in the expression of pride and shame is $0.16$ for Hollywood and $0.21$ for Bollywood.
Social expectations evoking shame are similar across movie industries (See Fig. 7 and 8). Sexuality (e.g., promiscuity, immodesty) is the dominant theme for sanctioning women in both movie industries whereas males are shamed for incompetency. Women across movie industries express more pride in family roles. Men-pride in Bollywood is centered on justice, winning and bravery whereas it is duty, self-identity and winning in Hollywood (See Fig. A1 for gender differences in pride).
Figure 7: Gender differences in shame-related themes in Bollywood movies. * indicates significant difference i.e., $p<0.05$.
Figure 8: Gender differences in shame-related themes in Hollywood movies. * indicates significant difference i.e., $p<0.05$.
@@ 5 Discussion
We release a cross-cultural dataset of shame- and pride-related expressions with implicit reasons behind them. To the best of our knowledge, the only related corpus is the GoEmotions dataset Demszky et al. (2020) comprising 817 samples for embarrassment and 452 samples for pride. Our dataset can help develop computational models for detecting shame/pride expressions and recognizing how appraisals of such emotions vary – enabling culturally cognisant human-computer conversations for diverse users Kim et al. (2011) and cross-cultural intent translation [^5] DARPA Computational Cultural Understanding: https://www.darpa.mil/program/computational-cultural-understanding.
@@
Shame is a highly undesirable emotion emphasizing incompetency and failures in the U.S. and used sparingly Cohen (2003); Boiger et al. (2013) whereas Eastern cultures use shame more frequently and it is known to have a desirable affect Lim (2016). Notably, we also observe that shame is 4.5 times more common in Bollywood compared to Hollywood. We also observe that shame is self-focused (reflecting internal shame) in Hollywood (See Table 2 and Fig. 2 for frequency distribution and psychosocial constructs). In contrast, shame is interdependent in collectivist communities Wong and Tsai (2007). The empirical analyses reveal other-focused (also, the public nature of shame) in Bollywood movies (See Fig. 3). Moreover, the contrasting tenses coupled with varying emotions, i.e., past-focus + sadness in Hollywood vs. present-focus + conflict in Bollywood, reflect their varying goals, i.e., remorse for past failures/losses vs enforcing conformity Wong and Tsai (2007). Pride-related discourse in Hollywood is duty and achievement-focused, in line with prior findings that the significance of "success" grows as a society becomes more individualistic Cohen (2003). In Bollywood, pride is collective and male-focused (we, achievement, male-focused) Khadilkar et al. (2022). It is also interesting to note linguistic markers such as determiner (used with objects/nouns) in Bollywood indicate more materialistic pride, compared to personal pronouns and social references in Hollywood which reflect more personal pride.
@@
Eliciting reasons leading to shame and pride expressions revealed social expectations across cultures. We observed high specificity in extracted reasons revealing cultural subtleties (e.g., desire for a son in Bollywood vs. returning calls after date night in Hollywood). To tease apart cultural differences, we first considered mapping social expectations/generated reasons to Schwartz’s Theory of Values Schwartz (2012). However, social expectations were found to have contrasting values depending on the culture. Consider "refusing to marry", an instance of non-conformity in Indian society, whereas it is an instance of self-direction in a Western context. We thus performed hierarchical clustering and empirically picked the distance after manually analyzing the quality and granularity of clusters. The situations connected to shame and pride in the U.S. society (i.e., lacking accountability, poverty, harm) vs Indian society (e.g., inappropriate sexual behavior, gender roles, betrayal - see Fig 4 and 5) align with known cultural tendencies (i.e., individualist vs collectivist) of both nations Triandis (1988).
@@
The vision of safe and accountable AI is centered on LLMs’ moral and value alignment. The dominant approach for norm discovery involves prompting LLMs, sometimes coupled with a verification step such as an entailment test or identifying underlying emotion (negative emotion $\rightarrow$ norm violation) Jiang et al. (2021); Fung et al. (2022); CH-Wang et al. (2023). However, a majority of social situations and the human annotators employed to label those situations reflect English beliefs and ethics. Our approach for extracting culture-specific social expectations and attitudes using social emotions overcomes this limitation. While the clustering was performed on the reasons extracted from movie conversations, the cluster themes are similar to patterns seen during LIWC analysis, validating our approach for overcoming Anglocentric bias in LLMs for norm discovery.
@@
Lastly, we examined how social expectations associated with shame/pride are attributed to men and women across cultures influencing their social behaviors. For example, women are less assertive across cultures to avoid negative attribution Amanatullah and Morris (2010); Ferguson and Eyre (2000) and ask for less during negotiation than men Arnold and McAuliffe (2021). We demonstrated that shame, a negative self-conscious emotion expressing devaluation, is targeted toward women more than men, and pride, a social emotion endorsing social value, is more used for men than women. The similarity in social expectations (e.g. regulating sexual behaviors of women vs shaming incompetency in men) across cultures is surprising. We thus feel that it is important to characterize social biases in the training data before their use for aligning LLMs.
@@ 6 Conclusion
We introduced a multi-cultural dataset of shame/pride-related expressions and the underlying social expectations. This study demonstrates (a) cross-cultural linguistic differences in shame and pride-related expression, offering insights into their functions across cultures, (b) cultural dichotomy in social expectations, and (c) more "social sanctions" and fewer "endorsements" to women in movies. Future work can utilize our dataset for culturally aligned LLMs and build social emotion perception and appraisal in human-computer interactions. Additionally, our analysis of shame/pride is the first of its kind; we hope future NLP researchers will build upon this work to investigate social expectations in LLMs from a multicultural, social emotion-based perspective.
@@
Social norms discovery is a crucial component of social and behavioral change programs [^6] ALIGN-https://www.alignplatform.org/learning-collaborative promoting equity, social justice, and well-being Mauduy et al. (2022); Bonan et al. (2020). Further work explores style as a product of norm differences Havaldar et al. (2023a). Social psychology investigates social norms (descriptive vs injunctive) to design experiments for understanding behaviors such as self-regulation, persuasion Cialdini et al. (1990) and decision-making Gavrilets (2020); Bhanot (2021) to promote collective-level change in societies.
@@
Social norms mutate as society evolves. We acknowledge that our dataset of movies (released post-1990) may reflect social norms that are less characteristic of contemporary society. Moreover, countries like India and America contain a mix of cultures. The captured norms may not reflect the cultural variations, for example, between regions (e.g., East Coast vs West Coast in the U.S.A or North vs South in India). Movies also exaggeratedly depict the world around us (e.g., weddings, criminal activities, sexual abuse, etc.), and we caution against stereotyping cultures based on movie-based norms.
The dominant language in Bollywood movies is Hindi and our analysis is based on their English translations which may not always be accurate, especially when the discourse is about concepts native to a culture. LIWC dictionary may also lack complete coverage for such concepts. We did not compare the movie genre and acknowledge that situational/unrealistic norms (e.g., science fiction, comedy, etc.) could exist. Regardless, social norms associated with shame and pride are still relatable and reflect the target audience’s beliefs. This study is conducted over a thirty-year long period and the aggregated norms may not truly reflect the current trends.
@@ Acknowledgements
This work was supported in part by the Penn Global Engagement Research Fund. We acknowledge the use of ChatGPT, an AI language model developed by OpenAI, for its assistance with writing.
@@ A.2 Examining Implicit shame
We prompted GPT-4 chat to identify (a) if a given discourse expresses shame, and (b) identify the behavior or action behind shame, for a randomly sampled set of 100 samples without explicit mentions of shame/pride each from Hollywood and Bollywood. We asked two human annotators to evaluate the correctness of GPT-4 chat-generated outputs. See Table A4 for annotation guidelines.
@@ A.3 Prompts
The prompts designed to extract the subject experiencing shame and pride, their gender, and the reason behind the emotions are provided in Tables A2 and A3.
@@
"You are an experienced social scientist and you study Indian society. Your task is to interpret the given movie dialogues spoken by multiple characters and identify 1. Who is being shamed? Predict the gender. If gender is not explicit, use the pronouns, name, and salutations to guess gender. 2. Identify the primary reason for shaming. If not explicit, return NA. The answer should be short and in CSV format. $<$ who_shame, gender, reason$>$ Input: And should we bow before others begging….them to marry our daughters? This shall not happen. Neither will the girls be alive here nor shall….we be ashamed of ourselves. You cannot kill the life which God has given. I won’t let you commit the sin. Output: girl’s parent, NA, not able to marry off their daughters Input: Black marketers are now in the open. And the thieves too Politics is in a great mess Shame on this system. There’s no democracy Get rid of these politicians The gong has struck..”Our hearts are swaying to it’s beats” Output: System, NA, poor law and regulations"
@@
"You are an experienced social scientist and you study American society. Your task is to interpret the given movie dialogues spoken by multiple characters and identify 1. Who is being shamed? Predict the gender. If gender is not explicit, use the pronouns, name, and salutations to guess gender. 2. Identify the primary reason for shaming. If not explicit, return NA. The answer should be short and in CSV format. $<$ who_shame, gender, reason$>$ Input: You still owe me 100. Remember? You stiffed Donny for 100 bucks? Cheapskate. Shame on you. Pay this man his C-note. Now I know why they call you the Snake. Output: NA, male, not returning borrowed money Input: You prey on your own people. You steal from your own people. Have you no shame!? - Huh? - Well, we’re still here. Man: Mr. Markopolos, it’s all yours. Output: Snake, male, stealing and preying on people"
@@ A.4 LIWC Correlation Results
Tables A7, A8, A9 and A10 contain positively correlated ($p<0.05$) LIWC categories, the most frequent five words for each category, Pearson r and confidence interval.
@@
$2^{nd}$ person pronouns (YOU) (you, your, you’re, yourself, you’ve) 0.161 0.000 [0.145, 0.178]
@@
Auxiliary verbs (AUXVERB) (is, are, have, be, don’t) 0.029 0.001 [0.013, 0.046]
@@
Table A7: Psychosocial categories positively correlated ($p<0.05$) with shame in Bollywood. p-values were corrected using Benjamini-Hochberg correction. The categories are arranged in decreasing order of correlation.
@@
Table A8: Psychosocial categories positively correlated ($p<0.05$) with pride in Bollywood. p-values were corrected using Benjamini-Hochberg correction. The categories are arranged in decreasing order of correlation.
@@
Table A9: Psychosocial categories positively correlated (p<0.05) with shame in Hollywood. p-values were corrected using Benjamini-Hochberg correction. The categories are arranged in decreasing order of correlation.
@@
Table A10: Psychosocial categories positively correlated (p<0.05) with pride in Hollywood. p-values were corrected using Benjamini-Hochberg correction. The categories are arranged in decreasing order of correlation.